Outstanding jobs from the ISIS action list

priority: 1-3

#12 
DATE: 15/5/93
PRIORITY: 3
Check efficiency of 2-nd order light with 1200 grating in blue
in relation to that of the 2400 holographic grating which is 
inefficient below 4000 A. Let Robin Clegg know.
STATUS: 
RESULT: -

#47
DATE: 30/8/93
PRIORITY: 2
STATUS: 
re-focussing the TV causes the image to change position. May be
that the focus drive is skew, or something is mis-alligned. One option
is that the TV mirror is misalligned after the quick-and-dirty job 3 
months ago. 
Also the TV focus is near one end of it's travel, which should be re-adjusted.
Contact Sue W. for good re-allignement job.
7/11/94 still not been solved; work ongoing.
RESULT: -

#51
DATE: 1/8/93
PRIORITY: 3
guiding is usually suspended while doing arc exposures
during observing. Maybe this can be improved by extra
baffling in the A&G
STATUS: 1/8/93: Sue looking into this
RESULT: -

#52
DATE: 26/10/93
PRIORITY: 2
whole of A&G box needs to be re-alligned by some optics expert !
STATUS: 8/11/93: talked with Sue Worswick who agrees. There should
be an opportunity during one of the stand-downs next year. See #47.
7/11/94: did not happen semester 94b; now trying to have it done
in semester 95a.
RESULT: -

#56
DATE: 9/11/93
PRIORITY: 3
get large quarter wave plate. There is not really a strong
scientific case for this since most objects for circular pol.
measurements are point sources anyway.
STATUS: -
RESULT: -

#57
DATE: 9/11/93
PRIORITY: 1
investigate possibility of improving response of Deuterium lamp,
or get any other kind of blue flat field lamp.
STATUS: 1/11/95: Dave King will do experiments to see what lamp might
produce good blue flats.
RESULT: -

#60
DATE: 19/12/93
PRIORITY: 2
measure which gratings have power by getting optimal focus
for all gratings. A systematic difference between gratings
will help keeping ISIS in optimal focus.
STATUS: 6/5/95: wait for ISIS servicing to be ready, then
re-establish *all* optimal focus values.
RESULT: -

#62
DATE: 15/2/94
PRIORITY: 2
blue grating positioning is off. 
STATUS: 15/2/94: improved, but still not what it used to be.
There now is an offset of +35 to the cenwave command to get the
correct wavelength centered (that is with the 300 grating).
RESULT: -

#63
DATE: 15/2/94
PRIORITY: 2
TV image shows flexure of a few (~2 arcseconds, ~10 pixels) which 
may be responsible for the somewhat poor calibrate results lately.
STATUS: 29/6/94: this has been thoroughly investigated and indeed
the TV probe was found to be flexing. This problem probably has
always been there. There is a plan to construct an end-stop
which would hold the probe in place to prevent is from flexing
26/1/95: still under investigation.
1/11/95: Cass A&G box has again thouroughly been looked at; the
TV probe has been tightened, but the effect on the sky now appears
worse than before. Continue investigating problem.
RESULT: -

#68
DATE: 12/4/94
PRIORITY: 2
ghosts have been reported with the new dichroics. Investigate what is
the cause.
STATUS: 12/4/94: images were taken with the narrow dekker and the
tungsten lamp. For fixed wavelengths in the read and the blue arm,
using the 600 gratings, spectra were taken through the different
dichroics and mirrors. Main results are listed elsewhere in detail.
Stray light is getting into the blue arm with the dichroic, most likely 
through reflections off the mirror next to the dichroic. Possibly a
baffle just above the dichroic slide which only admits light going
through the relevant path may reduce this problem.
This must be tested.
RESULT: -

#71
DATE: 11/5/94
PRIORITY: 2
have facility to store TV images
STATUS: 1/12/95: Stuart Barker looking into the general problem of
storing TV images, also for the purpose of measuring QE and S/N on
the TV systems for development of CCD based TV systems.
RESULT: -

#72
DATE: 11/5/94
PRIORITY: 2
have facility to store autoguider output
STATUS: 7/11/94: Bruce Gentles managed to get autoguider images
into IRAF format on the Sparc station. This has to be documented
and tested so we can release it as an option to observers.
RESULT: -

#73
DATE: 11/5/94
PRIORITY: 2
link GSS output to TCS and ICL so that probe coordinates don't
have to be typed in.
STATUS: 7/11/94: Robert Laing working on this. In pre-release version
it had been proven to work.
RESULT: -

#79
DATE: 1/6/94
PRIORITY: 2
We need a good ThAr lamp which is bright enough to be useful.
STATUS: 6/5/95: ThAr may have too many lines to be of general use
with ISIS. Investigate other arc lamps.
RESULT: -

#81
DATE: 12/8/94
PRIORITY: 2
it is found that the 6100 dichroic degrades the image quality in
the spatial direction. A narrow dekker (1.2") profile without
the dichroic results in a ~square spatial profile of ~4 pixels
wide, while through the dichroic the profile degrades to a
somewhat triangular shape with ~8 pixels base width. Note that
the degrading in the spectral direction was hardly noticable
(no focussing was attempted; collimator position was 5200, so
close to nominal). This was found at 8200 A with the R1200R 
grating and the EEV CCD on the red arm.
This needs to be checked at other wavelengths, with another grating,
and with the other dichroics.
STATUS:
RESULT:

#82
DATE: 10/11/94
PRIORITY: 2
Tom Marsh reports:
The blue arm has some really terrible
features in the flat field. Some of them are clearly internal reflections 
which do not affect the data and thus make the data worse. There are some 
other which are curved and run almost but not quite parallel to the slit.
These features are really bad, as much as 20% in fact, and for example
I think they have ruined the Hgamma absorption line in the data.
These may have been because no dekker was in place, but perhaps that 
should be made standard practice or something. Feel free to pass this onto
whoever is in charge of ISIS. The red arm flats have a very unpleasant
cusp. These really need to be cleared up I think,

the setup was 5700 dichroic, R600B grating, cenwave=4597

There is a most unpleasant reflection and unfortunately the object is
slap bang in the middle of it. there also appear to be 10 - 15 %
variations which run all across the chip. these things are all on hgamma.
worse, there were 2 nights of data with different formats and this
ultimately means that the hgamma cannot be used at all for radial velocities
as its breadth is very similar to the flat field features. 
STATUS: 6/5/95: this awaits finishing of the servicing and re-alignment
of ISIS.
RESULT: 15/1/95: the curved reflection Tom mentiones is a feature
which was introduced after the ISIS refurbishing last summer. Test so
far suggest that it is only there when the dome lights are on, and thus
should not affect the data when dome lights are off.
The other effects need more checking out.

#84
DATE: 25/1/95
PRIORITY: 2
we need red supression filters for both the red and blue arm of 
ISIS in order to get more balanced flat field with Tungsten lamp.
Also blue cutoff filters for the blue arm are required.
Finally an ND1.0 filter for blue and red arm would be useful.
STATUS: 25/1/95: ordered with Dave King:
for the RED arm of ISIS:
GG395, BG28, BG39 and ND1.0
size: 35.0 by 80.0 mm (tolerance 0.1 mm)
for the BLUE arm of ISIS:
BG28, ND1.0, GG495, RG630 and OG550
size: 27.9 by 80.0 mm (tolerance 0.1 mm)
RESULT:

#87
DATE: 22/2/95
PRIORITY: 1
acquire second small feed, silver coated, to replace the
large feed flat for WHIRCAM work on AUX port.
STATUS: 6/5/95: this is now a genuine RGO project. Push for
progress...
RESULT:

#88
DATE: 4/4/95
PRIORITY: 2
measure the actual spectral range convered for each grating,
and check whether it varies with wavelength. It appears that
the values listed in the manual are slightly too large.
STATUS:
RESULT:

#89
DATE: 4/4/95
PRIORITY: 2
update ISIS manuals and guide. Get things right regarding cross
disperser, fibres, polarimetry, dichroics, throughput, multi-masks,
etc.
STATUS:
RESULT:

#90
DATE: 4/4/95
PRIORITY: 2
update ISIS polarisation manual. Improve readibility and include
imaging polarimetry
STATUS:
RESULT:

#91
DATE: 6/5/95
PRIORITY: 1
improve the way the dichroic filters are mounted. Currently they
are held in place poorly by metal clamps which may work loose,
introducing serious flexure.
STATUS:
RESULT:

#95
DATE: 6/5/95
PRIORITY: 1
investigate possible improvements of ISIS structural flexure.
STATUS: 6/5/95: Spoken to Sue W about this. She thinks there is
scope for improving flexure. In the design phase the limits on
weight and off-axis moments for the Cass rotator were tight.
Now they are more relaxed and we could look into improving
the structural stiffness. This needs some careful mechanical
analysis.
RESULT:

#96
DATE: 15/1/96
PRIORITY: 1
Vik wrote: Every evening I check the focus with
isis_focus and every evening it changes by
up to 1000 collimator units! Having started the
run at the optimal collimator setting three nights 
ago I am not at very close to the upper limit of
the allowed range. Do you know why this might
be happening? I don't think it is due to people
touching the system since it is the weekend and 
the system has been left exactly as I left it.
What I have been doing every afternoon is taking
the system down and then bringing it back up again 
immediately. Perhaps this moves collimator slightly
(it never used to). I don't think the problem is
temperature fluctuations either. Very strange,
STATUS:
RESULT:

#97
DATE: 14/1/96
PRIORITY: 2
measure the actual polarisation through the polaroids in the
A&G box over the full spectral range, up to 1 micron.
Do this for the polaroids, but also for the calcite in the
main filter slide. Aim is to document the exact reange of which
the can be used as efficient polarizers
STATUS:
RESULT:

#99
DATE: 1/6/97
PRIORITY: 1
anecdotal reports on astigmatism of ISIS.
Check optimal collimator position
STATUS: 1/7/97: checked auto-collimation position on both arms.
Optimal values have not changed substantially since servicing
and realignment of ISIS in 1995.
RESULT: released

#100
DATE: 1/3/98
PRIORITY: 2
Measure the filter focus offsets for the rest of the ISIS colour 
filters in the red and blue arms. Antonio expressed some interest 
in carrying out these tests as a training programme. Also check if
focus position varies with grating. 
STATUS:
RESULT:


#103
DATE: 15/4/98
PRIORITY: 1
Investigate the Marsh suggestion that a fibre be used to direct light
from the A&G box onto the slit. Would effectively mean that arcs
can be done simultaneously with target exposures.
STATUS: 
RESULT:

#104
DATE: 20/2/98
PRIORITY: 2
Get a reflectivity test done on the ISIS dichroics and the mirrors
in each slide, from the comparisons there are some quite serious
deficiencies in the mirror reflectivities. 
STATUS: 
RESULT:


#104
DATE: 20/2/98
PRIORITY: 2
Get ISIS webpages updated, and try Peter Sorensons template. 
STATUS: Done
RESULT: ISIS webpages updated (20/2/2000) 


Javier Licandro (ISIS Instrument Specialist) licandro@ing.iac.es